OpenStudy (anonymous):

If you have a negative and a positive exponent..do you multiply, add, subtract, or divide

OpenStudy (anonymous):

i saw you multiply (:

OpenStudy (anonymous):

say*

hero (hero):

Negative exponent means "inverse" For example \(6^{-2}\) means inverse of \(6^2\)

OpenStudy (anonymous):

^?

hero (hero):

And the inverse of \(6^2\) is \(\huge \frac{1}{6^2}\)

OpenStudy (anonymous):

so that negative exponent is going to turn to a positive one and when did you see me multiply @ineedhelp1011111

OpenStudy (anonymous):

If you multiply the exponents of two coefficients, then you add them (unless one exponent is raised to another exponent - then you multiply them). Example: (x^3)(x^-2) = x^1 BUT (x^3)^-2 = x^-6

OpenStudy (anonymous):

thanks @richterp for the example

hero (hero):

@JazzyPoowh, negative exponent means "inverse". It is a symbol that indicates that the exponent is a fractional one. Negative exponent does not mean negative number.

hero (hero):

\(6^2 = 36\) however \(6^{-2} \ne -36\)

OpenStudy (anonymous):

this is my equation.......... The size of a certain cell is 2.5 × 10^–9 m. Another cell is 1.5 × 10^3 times larger. How large is the larger cell in scientific notation? (1 point)

OpenStudy (anonymous):

now do yhu see what im sayin @Hero

hero (hero):

Multiply them together

OpenStudy (coolshubhs):

if it is like,then: 5+-5 =5-5

hero (hero):

\(\frac{2.5}{10^9} \dot\ (1.5 \times 10^3)\)

OpenStudy (anonymous):

kk i multiplied them and got -27

hero (hero):

No, that's not what you get. You should review rules of exponents

OpenStudy (anonymous):

yhu just told me to multiply my exponents and that what i did -9 times 3 is equal to -27